OpenStudy (anonymous):

Which of the follow functions are linear? Explain how you know. y - 4 = -8(x – 1) y = x4 y – x2 = 4.5 3x + 5y = 15

OpenStudy (sparrow2):

in the second is it y=x*4 or y=\[y=x ^{4}\]

OpenStudy (sparrow2):

?

OpenStudy (anonymous):

y=x^4

OpenStudy (sparrow2):

then only first and last

OpenStudy (sparrow2):

linear mean that both x and y are in 1 degree

OpenStudy (anonymous):

same with the third x^2